OpenStudy (clamin):

HELP!! MEDAL!! Convert 3.5 meters/second to kilometers/hour. (please show step-by-step)

OpenStudy (johnweldon1993):

\[\large 3.5 \frac{m}{s}\] We know there are 1000m in 1 km so \[\large 3.5 \frac{\cancel{m}}{s} \times \frac{1km}{1000\cancel{m}}\] Now notice the 'm' will cancel So and now we also know there are 3600 seconds in 1 hour...so \[\large 3.5\frac{\cancel{m}}{\cancel{s}} \times \frac{1km}{1000\cancel{m}} \times \frac{3600\cancel{s}}{1h}\] So finally we have \[\large (3.5 \times \frac{1}{1000} \times 3600)\frac{km}{h}\]